OpenStudy (anonymous):

let g(x) be a horizontal compression of f(x)=3x+5 by a factor of 1/2. Write the rule for g(x) and graph the function.

OpenStudy (anonymous):

4 options: g(x)=2x+5 g(x)=3/2x+5 g(x)=3/2x+5/2 g(x)=6x+5

OpenStudy (anonymous):

so D?